1) how many different ways can the letters in the word STATISTICS be arranged if all the letters T are seperated?
2) a group of five people is to be formed from 6 men and 8 women for a trip.In how many ways can the five people be selected if John and Mary must go on the trip and there are more men than women?

Answer by stanbon(75887)

1) how many different ways can the letters in the word STATISTICS be arranged if all the letters T are separated?
------
If you treat the the 3 T's as one letter there are only 8 letters
to arrange:
Ans: 8!/(3!*2!) = 3360 arrangements
----------------------------------------------
2) a group of five people is to be formed from 6 men and 8 women for a trip.
In how many ways can the five people be selected if John and Mary must go on the trip and there are more men than women?
-----
Acceptable mix: 5m/3w ; 6m/2w
# of 5/3 groups with J and M included = 5C4*7C2 = 5*21 = 105
---
# of 6/2 groups with J and M included = 5C5*7C1 = 7
---
Cheers,
Stan H.